Problem #1 Calculate Z, HR, and SR by the Redlich/Kwong equation for each of the following substances, and compare the results with values found from suitable generalized correlations: a. Benzene at 575 K and 30 bar. b. Ethylene at 300 K and 35 bar. Problem #2 Calculate Z, HR, and SR for benzene at 575K and 20 bar. You can assume that the pressure is low enough to apply the correlation for the second virial coefficient. Problem #3 In a(T)I In(T,) Prove the equation in the red box for each one of the EOS.
Answer #1

SOK ㄧㄚ 126:2K 3 3.9 1g 1-S a . O0866 3.8IOTえ0-1095 = (270-1o2, 굿)帐 (-1095-04096굿 え= 0.6611 ヌ七e0.66リン 1So RT RT 666t PT0 s536) 3869 - 898
Answer #2

EX1: Calculate Z, HR, and SR by the Redlich/Kwong equation for Benzene at 575 K and 30 bar, and compare result with value found from suitable generalized correlation.